# Referee: Reference-aware Audiovisual Deepfake Detection

### v1

@@ Abstract
Since deepfakes generated by advanced generative models have rapidly posed serious threats, existing audiovisual deepfake detection approaches struggle to generalize to unseen forgeries. We propose a novel reference-aware audiovisual deepfake detection method, called Referee. Speaker-specific cues from only one-shot examples are leveraged to detect manipulations beyond spatiotemporal artifacts. By matching and aligning identity-related queries from reference and target content into cross-modal features, Referee jointly reasons about audiovisual synchrony and identity consistency. Extensive experiments on FakeAVCeleb, FaceForensics++, and KoDF demonstrate that Referee achieves state-of-the-art performance on cross-dataset and cross-language evaluation protocols. Experimental results highlight the importance of cross-modal identity verification for future deepfake detection. The code is available at https://github.com/ewha-mmai/referee.
Index Terms—  Audiovisual deepfake detection, speaker verification
@@ 1 Introduction
With the rapid development of generative AI, recent models are increasingly capable of synthesizing media content, including both video and audio. However, this advancement carries a dual nature: while it enables creative and beneficial applications, it also introduces significant risks of misuse and societal harm. The risk has arisen from synthesized media that convincingly imitate humans, known as deepfake [1].
The goal of deepfake detection is to distinguish genuine (i.e., real) or spoof (i.e., fake) content. Prior works have relied on a single modality. Image-based approaches have focused on detecting shape distortion [2, 3], boundary artifacts [4, 5, 6], or anomalies in the frequency domain [7, 8] that arise during the synthesis process. On the other hand, audio-based approaches [9, 10] have learned discriminative features to capture vulnerable spectral patterns. Although these methods have shown promising performance, the synthetic contents are generated with higher fidelity, making unimodal detection strategies less reliable. On the other hand, audiovisual deepfake detection [11, 12, 13] has exploited the natural correspondence between facial movements and speech. The dynamic motion patterns [14, 15, 16], such as lip synchronization, or unnatural temporal transitions, have been investigated to capture manipulation cues.
However, such methods often fail when the lips are partially occluded by hands, microphones, or head poses, and their performance further degrades under low-resolution conditions.
Fig. 1: Referee introduces robust audiovisual deepfake detection guided by a one-shot reference example, verifying the cross-modal biometrics consistency as well as temporal unnaturalness.
When humans encounter a video, they do not fixate on lip motion in isolation; instead, they weave together subtle cues, e.g., tone of voice, facial appearances, and identity coherence, to form a judgment of authenticity. Motivated by this natural ability, we introduce Referee, a novel model that examines the consistency of speakers’ identity for robust audiovisual deepfake detection. In contrast, artifact-based or synchronization-based methods often rely heavily on the generative models used to create deepfakes in the training set, which limits their generalizability to unseen forgeries [17, 18]. As generative techniques continue to advance, such low-level visual or acoustic artifacts become increasingly subtle or even imperceptible. By comparison, preserving a speaker’s unique identity across modalities remains a highly challenging task for generative models. While recent work [19] has proposed a method to measure speaker consistency using contrastive learning, it requires long reference footage (i.e., at least 10 samples of 30 seconds per person), which is often impractical in real-world scenarios.
Fig. 2: The overall framework of Referee. Target and reference videos are encoded into audiovisual features, which are passed through IDB module to generate identity queries. The identity matching module refines the target tokens with respect to the reference tokens, after which the reference-aware identity queries, together with the target audiovisual features and a [CLS] token, are processed by the AV-Transformer for deepfake classification and identity matching.
@@
Referee presents an identity bottleneck module (IDB) with learnable identity queries to encode audiovisual sequences into a unified identity representative feature, as shown in Fig. 1. Identity queries obtained from both reference and target content are subsequently aligned through cross-attention to verify speaker identity. This process is optimized by our ID matching loss to ensure that the models can capture speaker-specific cues rather than superficial artifacts. The fused identity representation is then attached to the audiovisual feature, enabling the system to reason jointly about two critical aspects: (1) whether lip-speech synchronization is consistent over time, and (2) whether either or both modalities have been manipulated in a way that violates identity integrity. Finally, the model integrates these signals to classify the content as real or fake. Extensive experiments show Referee effectively surpasses current state-of-the-art works on FakeAVCeleb [20], FaceForensics++ [21], and KoDF [22]. It also demonstrates that Referee is inherently more resilient to the rapid progress of generative models and better equipped to handle unseen types of deepfakes.
To sum up, our contributions are:
@@
  • Extensive experiments on FakeAVCeleb, FaceForesics++, KoDF demonstrate that our Referee achieves state-of-the-art performance on both unseen forgery manipulations and cross-lingual evaluation.
@@ 2.1 Motivation and overview
Most prior approaches [11, 12, 13] for audiovisual deepfake detection have shown impressive performance while concentrating on correct synchronization between audio and video streams. However, there are several challenges: (1) methods relying on synchronization are inherently tied to lip-sync manipulation methodologies, making it difficult to generalize to unseen forgeries. (2) These methods also remain vulnerable under occlusion, low resolution, or degraded signals. To solve these limitations, we propose Referee to detect forgery with speaker identity information. We claim that cross-modal biometric cues [23, 24], such as verifying the correspondence between the speaker’s voice and facial appearance, provide a strong basis for deepfake detection.
@@ 2.2 Audiovisual feature extraction
Given an input video and audio, our video and audio encoders extract visual and audio features, respectively. In this work, pretrained audiovisual encoders [25] are used to initialize the parameters. The video and audio are divided into $N_{seg}$ fixed-length segments of length $T$ seconds. From each of the $N_{seg}$ segments, the visual and audio encoders yield segment-level token embeddings. To construct unified representations, these embeddings are concatenated along the temporal axis, forming continuous token sequences for each modality. Formally, this process yields the following visual and audio feature sequences:
$F_{v}^{*}\in\mathbb{R}^{(N_{\text{seg}}\cdot T_{v})\times D},$ (1)
$F_{a}^{*}\in\mathbb{R}^{(N_{\text{seg}}\cdot T_{a})\times D},$ (2)
where * denotes either the target (TGT) or reference (REF), $T_{v}$ and $T_{a}$ denote the number of visual and audio tokens per segment, and $D$ is the feature dimension.
Finally, for both inputs, the visual and audio sequences are concatenated to form joint audiovisual representations. These are formally expressed as:
$F^{\text{TGT}}$ $=[F_{v}^{\text{TGT}};\,F_{\text{mod}};\,F_{a}^{\text{TGT}}],$ (3)
$F^{\text{REF}}$ $=[F_{v}^{\text{REF}};\,F_{\text{mod}};\,F_{a}^{\text{REF}}],$ (4)
where $F_{\text{mod}}$ is a special learnable token separating the two modalities. These two joint representations, $F^{\text{TGT}}$ and $F^{\text{REF}}$, serve as the input for subsequent identity reasoning and classification modules.
@@ 2.3 Identity bottleneck (IDB)
The entanglement of speaker identity with factors like speech content, facial expression, and pose within audiovisual features presents a significant challenge for robust identity (ID) comparison. To address this, we introduce a new ID bottleneck (IDB) that captures ID-related features into a core identity representation.
Inspired by prior works [26, 27], our IDB employs a set of learnable queries to extract and summarize cross-modal identity cues. The module consists of $L$ identical blocks, each composed of a self-attention layer, a cross-attention layer, and a feed-forward network. The process within each block begins with a set of $N_{q}$ learnable identity queries, $Q_{(0)}\in\mathbb{R}^{N_{q}\times D}$. To stabilize learning and encode relative positions, each query is initialized with a learnable positional embedding. For the $l$-th block $(l=1,\dots,L)$, the queries are updated in three stages. At each block, the queries first undergo self-attention, allowing them to capture internal dependencies and integrate information among themselves. The resulting intermediate queries, $\tilde{Q}_{(l)}$, then act as the query input to a multi-head cross-attention layer, where the joint representation ($F^{\text{TGT}}$ or $F^{\text{REF}}$) serves as the key and value. The output of the cross-attention layer is passed through a feed-forward network to produce the block’s final output, $Q_{(l)}$. This three-stage update process is formally defined as:
$\tilde{Q}_{(l)}$ $=\texttt{SA}\big(\texttt{LN}(Q_{(l-1)})\big)+Q_{(l-1)},$ (5)
$\hat{Q}_{(l)}$ $=\texttt{CA}\big(\texttt{LN}(\tilde{Q}_{(l)}),\texttt{LN}(F)\big)+\tilde{Q}_{(l)},$ (6)
$Q_{(l)}$ $=\texttt{FFN}\big(\texttt{LN}(\hat{Q}_{(l)})\big)+\hat{Q}_{(l)}$ (7)
Repeating this three-stage process over $L$ blocks progressively refines the queries into tokens that encode essential identity information. This yields two distinct sets of identity tokens: $T^{\text{TGT}}_{\text{ID}}=Q_{(L)}^{\text{TGT}}=\texttt{IDB}(F^{\text{TGT}})$ from the target features, and $T^{\text{REF}}_{\text{ID}}=Q_{(L)}^{\text{REF}}=\texttt{IDB}(F^{\text{REF}})$ from the reference features.
2.4 Deepfake detection
@@
Referee performs deepfake detection by jointly leveraging fine-grained audiovisual correspondence cues and identity consistency information between the target and the reference. Our ID matching module aggregates and compares identity consistency, which employs $M$ stacked cross-attention blocks. The target tokens serve as the query, while the reference tokens act as the key and value. Namely, the target query $T^{\text{TGT}}_{\text{ID}}$ is refined by conditioning it on the $T^{\text{REF}}_{\text{ID}}$. This matching process generates reference-aware identity tokens, $\bar{T}^{\text{TGT}}_{\text{ID}}$, which not only retain the intrinsic identity characteristics of the target but also embed information about similarities and discrepancies with the reference. As a result, even subtle inconsistencies between the two identities can be amplified and clearly manifested in the final representation.
Subsequently, $\bar{T}^{\text{TGT}}_{\text{ID}}$ are concatenated with the [CLS] token and the target audiovisual features $F^{\text{TGT}}$:
@@
The resulting sequence $X^{\text{TGT}}$ is passed through joint audiovisual transformer blocks (shortly, AV-Transformer), which integrate temporal audiovisual factors (e.g., artifacts, synchronization cues) with identity-centric cues. To better distinguish different token roles, we incorporate type embeddings that explicitly encode whether a token corresponds to the classification token or identity queries. This role-specific bias encourages the model to facilitate distinct representations following their roles. Finally, a deepfake classifier takes the [CLS] token, which produces the final deepfake prediction.
To ensure the ID matching module produces highly discriminative features for identity consistency check, we design an auxiliary identity verification task. Specifically, the reference-aware identity tokens $\bar{T}_{\text{ID}}^{\text{TGT}}$ are aggregated via average pooling to form a single feature vector. This vector is then fed into a dedicated MLP head to predict the identity matching logit, $Y_{\text{ID}}$. This auxiliary objective encourages the matching module to learn features that are not only useful for the final deepfake detection task but also explicitly optimized for robust identity verification.
Our overall training objective is composed of a deepfake classification loss $\mathcal{L}_{\text{RF}}$ and an identity matching loss $\mathcal{L}_{\text{ID}}$, both implemented as a standard cross-entropy loss. The identity matching task verifies whether two videos contain the same identity. Regardless of whether the manipulation originates from the reference speaker, we treat the forged content as representing another identity. Equivalently, pairs were labeled as the same identity only if they came from distinct real recordings of the same individual.
By jointly learning the deepfake classification and the identity matching tasks, the model not only learns to detect whether an input video is manipulated but also acquires the ability to discriminate identities. This identity discrimination capability, in turn, enhances the effectiveness of deepfake detection.
@@ 3 Experiments
Method Modality ACC AUC
Xception [21] V 67.9 70.5
LipForensics [28] V 80.1 82.4
FTCN [29] V 64.9 84.0
CViT [30] V 69.7 71.8
RealForensics [31] V 89.9 94.6
VFD [32] AV 81.5 86.1
AVoiD-DF [16] AV 83.7 89.2
AVFF [14] AV 98.6 99.1
Referee AV 99.24 99.71
Method Train set Modality AUC AP
LipForensics [28] FF++ V 93.25 -
RealForensics [31] FF++/AVLips V 91.32 -
@@
Referee FakeAVCeleb AV 68.44 87.10
@@ 3.1 Datasets
FakeAVCeleb. We train Referee on FakeAVCeleb [20], which includes both audio and visual manipulations. Visual manipulations were created using FaceSwap [34], FSGAN [35], and Wav2Lip [36], while audio manipulations were generated with SV2TTS [37]. The dataset provides deepfake videos with only visual manipulation, only audio manipulation, or manipulations in both modalities. We further sourced 37k real videos corresponding to the same 500 identities from VoxCeleb2 [38] to use as reference videos.
FaceForensics++ (FF++). We use FF++ [21] to evaluate the generalization capabilities of methods across the difference of training and test data. The benchmark encompasses five manipulation methods: DeepFakes [39], FaceSwap [34], Face2Face [40], NeuralTexture [41], and FaceShift [42]. It includes data subjected to not only face swapping but also expression manipulation, enabling the evaluation of the models’ performance across diverse manipulation types. In the official test split, we only use the videos from which the corresponding audio tracks could be reliably extracted.
KoDF. To evaluate the performance on zero-shot cross-lingual forgery detection, we utilize KoDF [22], a large-scale deepfake dataset based on Korean speech generated using six different techniques: FaceSwap [34], DeepFaceLab [43], FSGAN [35], FOMM [44], ATFHP [45], and Wav2Lip [36]. Following [33, 14], we randomly sampled 100 real and 100 audio-driven manipulated fake videos for evaluation, except manipulation techniques already employed in FakeAVCeleb.
3.2 Implementation details and evaluation setup

### v2

# Boo Lee Lee

@@ Abstract
Deepfakes generated by advanced generative models have rapidly posed serious threats, yet existing audiovisual deepfake detection approaches struggle to generalize to unseen manipulation methods. To address this, we propose a novel reference-aware audiovisual deepfake detection method, called Referee, to capture fine-grained identity discrepancies. Unlike existing methods that overfit to transient spatiotemporal artifacts, Referee employs identity bottleneck and matching modules to model the relational consistency of speaker-specific cues captured by a single one-shot example as a biometric anchor. Extensive experiments on FakeAVCeleb, FaceForensics++, and KoDF demonstrate that Referee achieves state-of-the-art results on cross-dataset and cross-language evaluation protocols, including a 99.4% AUC on KoDF. These results highlight that explicitly correlating reference-based biometric priors is a key frontier for achieving generalized and reliable audiovisual forensics.
keywords:
deepfake detection, speaker verification
@@ 1 Introduction
Rapid advances in generative AI allow models [kowalski2018faceswap, jia2018transfer, flux2024] to synthesize high-fidelity multimodal content, including video and audio. Despite its creative potential, this capability poses significant risks of misuse and societal harm, exemplified by deepfake [tolosana2020deepfakes] that convincingly imitate humans. Deepfake detection aims to distinguish genuine (i.e., real) or spoof (i.e., fake) content. Prior unimodal approaches have primarily focused on image-based artifacts, including shape distortion [guo2022eyes, li2018exposing], boundary artifacts [li2020face, chen2023watching, fu2025faces], or frequency anomalies [tan2024frequency, li2021frequency]. Alternatively, audio-based methods [qais2022deepfake, chakravarty2024lightweight] have targeted discriminative spectral patterns to identify synthetic speech. Although these methods have shown promising performance, the synthetic content is generated with higher fidelity, making unimodal detection strategies less reliable.
In contrast, audiovisual deepfake detection [agarwal2020detecting, liu2024lips, zhou2021joint, oorloff2024avff] has exploited the natural correspondence between facial movements and speech. These methods have investigated dynamic motion patterns [liang2024speechforensics, yang2023avoid], such as lip synchronization and temporal transitions, to capture manipulation cues. However, because these approaches rely on fine-grained visual cues within the lip area, they remain vulnerable to partial lip occlusions caused by hands, microphones, or head poses. Furthermore, their effectiveness diminishes significantly in low-resolution conditions where such subtle spatial details are heavily degraded.
When evaluating video authenticity, humans do not fixate on lip motion in isolation; instead, they holistically integrate diverse cues, e.g., tone of voice, facial appearance, and identity coherence. Prior identity-based methods like ID-Reveal [idreveal] and ICT-Ref [ictref] rely on visual-only comparisons, often employing proxy tasks with heavy generative models or simple mask-based transformation. Such reliance can cause these models to overfit to low-level synthesis artifacts rather than capturing identity features. POI-Forensics [poiforensics] incorporates audiovisual data with naïve feature-level concatenation for contrastive learning, which precludes meaningful crossmodal exchange. This structural limitation not only prevents the model from capturing intricate cross-modal dependencies but also necessitates excessively long reference footage (e.g., over five minutes per identity), rendering it impractical for real-world deployment.
Figure 1: The overall framework of Referee. Identity queries are extracted by the IDB and conditionally refined in the ID matching module relative to the reference. These reference-aware queries, alongside target features, are processed by the AV-Transformer for deepfake classification, while the average-pooled reference-aware queries are used for identity classification.
@@
In this paper, we introduce Referee, a novel framework that jointly captures speaker identity inconsistency and audiovisual desynchronization for robust deepfake detection. Referee presents an identity bottleneck module (IDB) with learnable identity queries to encode audiovisual sequences into a unified identity representative feature. Identity queries obtained from both reference and target content are subsequently aligned through cross-attention to verify speaker identity. This process is optimized by our ID matching loss to ensure that the models can capture speaker-specific cues rather than superficial artifacts. The fused identity representation is then attached to the audiovisual feature, enabling the system to reason jointly about two critical aspects: (1) whether lip-speech synchronization is consistent over time, and (2) whether either or both modalities have been manipulated in a way that violates identity integrity. Finally, the model integrates these signals to classify the content as real or fake. Extensive experiments show Referee effectively surpasses current state-of-the-art works [oorloff2024avff, nie2024frade, anshul2025intra] on FakeAVCeleb [khalid2021fakeavceleb], FaceForensics++ [rossler2019faceforensics++], and KoDF [kwon2021kodf]. It also demonstrates that Referee is inherently more resilient to the rapid progress of generative models and better equipped to handle unseen types of deepfakes. To sum up, our contributions are:
@@
  • Extensive experiments on FakeAVCeleb, FaceForensics++, KoDF demonstrate that our Referee achieves state-of-the-art performance on both unseen forgery manipulations and cross-lingual evaluation.
@@ 2.1 Motivation and overview
Most prior approaches [agarwal2020detecting, liu2024lips, zhou2021joint, oorloff2024avff] for audiovisual deepfake detection have shown impressive performance while concentrating on correct synchronization between audio and video streams. However, there are several challenges: (1) methods relying on synchronization are inherently tied to lip-sync manipulation methodologies, making it difficult to generalize to unseen forgeries. (2) These methods also remain vulnerable under occlusion, low resolution, or degraded signals. To solve these limitations, we propose Referee to detect forgery with speaker identity information. We claim that cross-modal biometric cues [nagrani2018seeing, lee2023imaginary], such as verifying the correspondence between the speaker's voice and facial appearance, provide a strong basis for deepfake detection.
@@ 2.2 Audiovisual feature extraction
Given an input video and its corresponding audio, our framework first extracts modality-specific representations. To leverage robust temporal alignment, our visual and audio encoders are built on pretrained audiovisual encoders [iashin2024synchformer]. The input streams are divided into $N_{seg}$ fixed-length segments, each with a duration of $T$ seconds. For each segment, the encoders yield segment-level token embeddings, $F_{v}\in\mathbb{R}^{(N_{\text{seg}}\cdot T_{v})\times D}$ and $F_{a}\in\mathbb{R}^{(N_{\text{seg}}\cdot T_{a})\times D}$, where $T_{v}$ and $T_{a}$ denote the number of visual and audio tokens per segment, and $D$ is the feature dimension. To facilitate cross-modal interaction, the visual and audio sequences are concatenated to form a joint representation for both the target and reference samples. With a learnable modality-separation token $F_{\text{mod}}$, the joint audiovisual sequences are constructed:
$F^{\text{TGT}}=[F_{v}^{\text{TGT}};\,F_{\text{mod}};\,F_{a}^{\text{TGT}}],F^{\text{REF}}=[F_{v}^{\text{REF}};\,F_{\text{mod}};\,F_{a}^{\text{REF}}].$ (1)
@@ 2.3 Identity bottleneck (IDB)
A fundamental challenge in robust identity (ID) compression is that speaker identity is often inherently entangled with transient factors like speech content, facial expression, and head poses. To isolate identity-relevant signals from these confounding variables, we introduce a new ID bottleneck (IDB). Our IDB employs a set of $N_{q}$ learnable queries [jaegle2021perceiver, li2023blip], $Q_{\text{ID}}\in\mathbb{R}^{N_{q}\times D}$, to extract and summarize cross-modal identity cues. Crucially, we employ a weight-sharing scheme for the IDB across both target and reference branches. The module consists of $L$ identical transformer blocks, each composed of a self-attention layer, a cross-attention layer, and a feed-forward network. To stabilize learning and encode relative positions, each query is initialized with a learnable positional embedding.
Based on the cross-attention mechanism, the learnable queries ($Q_{\text{ID}}$) interact with the joint audiovisual feature sequences ($F^{*}$) in each transformer block, where the latter serve as the Key ($K$) and Value ($V$). This yields two distinct sets of identity tokens: $T^{\text{TGT}}_{\text{ID}}=\texttt{IDB}(Q_{\text{ID}},F^{\text{TGT}})$ from the target features, and $T^{\text{REF}}_{\text{ID}}=\texttt{IDB}(Q_{\text{ID}},F^{\text{REF}})$ from the reference features. Through this process, the queries adaptively attend to identity-salient regions within the audiovisual sequence while ignoring transient factors.
2.4 Matching cross-modal biometrics
@@
Our ID matching module is designed to evaluate cross-modal biometrics by quantifying the consistency between the target and reference samples. While the preceding IDB extracts individual biometric signatures, this module performs a cross-sample alignment to detect identity mismatches. We employ $M$ stacked cross-attention blocks where the target tokens (i.e., $T^{\text{TGT}}_{\text{ID}}$) serve as the Query, and the reference tokens (i.e., $T^{\text{REF}}_{\text{ID}}$) act as the Key and Value. This matching process generates reference-aware identity tokens, $\bar{T}^{\text{TGT}}_{\text{ID}}$, which not only retain the intrinsic identity characteristics of the target but also embed information about similarities and discrepancies with the reference. As a result, even subtle inconsistencies between the two identities can be amplified and clearly manifested in the final representation.
To ensure the ID matching module prioritizes identity-relevant cues, we design an auxiliary identity verification task to discern whether the target and reference samples originate from the same individual. Specifically, the reference-aware identity tokens $\bar{T}_{\text{ID}}^{\text{TGT}}$ are transformed into a global identity feature via average pooling, which is then fed into an MLP head to predict the identity matching logit, $\hat{Y}_{\text{ID}}$. This process is supervised by an identity matching loss $\mathcal{L}_{\text{ID}}$, implemented as a binary cross-entropy between the prediction and the ground-truth label $Y_{\text{ID}}\in\{0,1\}$. Crucially, any forged content, regardless of its intended target, is treated as a distinct identity from the reference speaker. Therefore, a pair is assigned $Y_{\text{ID}}$=1 exclusively when both samples originate from distinct, authentic recordings of the same individual. This auxiliary objective encourages the matching module to learn features that are not only conducive to the final deepfake detection task but also explicitly optimized for robust cross-modal identity verification.
2.5 Audiovisual deepfake detection
Referee performs deepfake detection by jointly leveraging fine-grained audiovisual correspondence cues and identity consistency information between the target and the reference. Subsequently, $\bar{T}^{\text{TGT}}_{\text{ID}}$ are concatenated with the [CLS] token $T_{\text{[CLS]}}$ and the target audiovisual features $F^{\text{TGT}}$:
@@ 2.5 Audiovisual deepfake detection
The resulting sequence $X^{\text{TGT}}$ is passed through joint audiovisual transformer blocks (shortly, AV-Transformer), which integrate temporal audiovisual factors (e.g., artifacts, synchronization cues) with identity-centric cues. Finally, the refined representation of the [CLS] token from the last transformer layer is fed into a deepfake classifier to produce the final deepfake prediction.
The final training objective, $\mathcal{L}_{\text{all}}$, is composed of a deepfake classification loss $\mathcal{L}_{\text{RF}}$ and the identity matching loss $\mathcal{L}_{\text{ID}}$:
$\mathcal{L}_{\text{all}}=\mathcal{L}_{\text{RF}}+\mathcal{L}_{\text{ID}},$ (3)
where $\mathcal{L}_{\text{RF}}$ denotes the standard cross-entropy loss for distinguishing real from forged content. By jointly optimizing these tasks, the model not only learns to detect generic manipulation artifacts but also develops a robust identity-discriminative capability, which in turn serves as a powerful prior for high-fidelity deepfake detection.
@@ 3 Experiments
Method Modality w/ Ref. Train set AUC AP
@@
[…]
Referee AV FakeAVCeleb 79.78 91.00
Table 2: Cross-dataset performance comparison on KoDF. $\ddagger$ denotes results evaluated using publicly available pretrained weights.
Method Modality w/ Ref. Train set AUC AP
Xception [rossler2019faceforensics++] V FakeAVCeleb 77.7 76.9
LipForensics [haliassos2021lips] V FakeAVCeleb 86.6 89.5
FTCN [zheng2021exploring] V FakeAVCeleb 68.1 66.8
[…]
RealForensics [haliassos2022leveraging] V FakeAVCeleb 93.6 95.7
AV-DFD [zhou2021joint] AV FakeAVCeleb 82.1 79.6
AVAD [feng2023self] AV LRS2/LRS3 86.9 87.6
[…]
AVFF [oorloff2024avff] AV FakeAVCeleb 95.5 93.1
FRADE [nie2024frade] AV FakeAVCeleb 92.4 -
ICSAV [anshul2025intra] AV LRS2 99.2 98.6
FoVB [nie2025towards] AV FakeAVCeleb 94.3 -
Referee AV FakeAVCeleb 99.4 99.4
@@ 3.1 Datasets
We train Referee on FakeAVCeleb [khalid2021fakeavceleb], which includes both audio and visual manipulations. Visual manipulations were created using FaceSwap [kowalski2018faceswap], FSGAN [nirkin2019fsgan], and Wav2Lip [prajwal2020lip], while audio manipulations were generated with SV2TTS [jia2018transfer]. The dataset provides deepfake videos with only visual manipulation, only audio manipulation, or manipulations in both modalities. We further sourced 37k real videos corresponding to the same 500 identities from VoxCeleb2 [voxceleb2] to use as reference videos. Following prior works [yang2023avoid, oorloff2024avff], FakeAVCeleb is split 70% for training, and 30% for a test set. To assess cross-dataset generalization, we evaluate on FaceForensics++ (FF++) [rossler2019faceforensics++], encompassing five manipulation methods [faceswap2018, kowalski2018faceswap, thies2016face2face, thies2019deferred, li2019faceshifter] ranging from face-swapping to expression manipulation. Evaluation is conducted on the official test split for videos where audio tracks could be reliably extracted. For zero-shot cross-lingual evaluation, we utilize KoDF [kwon2021kodf], a Korean speech deepfake dataset manipulated with six different techniques [kowalski2018faceswap, perov2020deepfacelab, nirkin2019fsgan, siarohin2019first, yi2020audio, prajwal2020lip]. Following [feng2023self, oorloff2024avff], we randomly sampled 100 real and 100 audio-driven manipulated fake videos for evaluation, except manipulation techniques already employed in FakeAVCeleb.
3.2 Implementation details